# DenoiseReID: Denoising Model for Representation Learning of Person Re-Identification

### v1

@@ Abstract
The denoising model has been proven a powerful generative model but has little exploration of discriminative tasks. Representation learning is important in discriminative tasks, which is defined as "learning representations (or features) of the data that make it easier to extract useful information when building classifiers or other predictors" in [4]. In this paper, we propose a novel Denoising Model for Representation Learning and take Person Re-Identification (ReID) as a benchmark task, named DenoiseReID, to improve feature discriminative with joint feature extraction and denoising. We first view each embedding layer in a backbone as a denoising layer, processing the cascaded embedding layers as if we are recursively denoise features step-by-step. This unifies the frameworks of feature extraction and feature denoising, where the former progressively embeds features from low-level to high-level, and the latter recursively denoises features step-by-step. Then we design a novel Feature Extraction and Feature Denoising Fusion Algorithm (FEFDFA) and theoretically demonstrate its equivalence before and after fusion. FEFDFA merges parameters of the denoising layers into existing embedding layers, thus making feature denoising computation-free. This is a label-free algorithm to incrementally improve feature also complementary to the label if available. Experimental results on 4 ReID datasets and various of backbones show the stability and impressive improvements. We also extend the proposed method to large-scale (ImageNet) and fine-grained (CUB200, Oxford-Pet, Flowers) classification tasks, similar improvements are observed. The code will be released on github.
@@ 1 Introduction
Denoising Diffusion Probabilistic Models (DDPM)  [20] or Diffusion Model for short have been proven to be a powerful generative model [5]. Generative models can generate vivid samples (such as images, audio and video) by modelling the joint distribution of the data $P(X,Y)$, where $X$ is the sample and $Y$ is the condition. Diffusion models achieve this goal by adding Gaussian noise to the data and training a denoising model of inversion to predict the noise. Diffusion models can generate multi-formity and rich samples, such as Stable diffusion [45], DALL [43] series and Midjourney, these powerful image generation models, which are essentially diffusion models.
However, its application to discriminative models has not been extensively explored. Different from generative models, discriminative models predict data labels by modelling the marginal distribution of the data $P(Y|X)$. $Y$ can be various labels, such as image tags for classification, object boxes for detection, and pixel tags for segmentation. Currently, there are several methods based on diffusion models implemented in specific fields. For example, DiffusionDet [6] is a new object detection framework that models object detection as a denoising diffusion process from noise boxes to object boxes. It describes object detection as a generative denoising process and performs well compared to previous mature object detectors. DiffSeg [47] for image segmentation, which is a method of unsupervised zero-shot sample segmentation using pre-trained models (stable diffusion). It introduces a simple and effective iterative merging process to measure the attention maps between KL divergence and merge them into an effective segmentation mask. The proposed method does not require any training or language dependency to extract the quality segmentation of any image.
Figure 1: A brief description of our idea. (a) A typical denoising model for generative tasks recursively applies a denoising layer. (b) A naive idea that applies a denoising strategy to a discriminative model is applying a recursive denoise layer on the feature of a backbone and taking extra inference latency. (c,d) Our DenoiseReID first unifies the frameworks of feature extraction and denoising in a backbone pipeline, then merges parameters of the denoising layers into embedding layers, making the feature more discriminative without extra latency cost.
@@
To resolve all the difficulties above and efficiently apply the denoising process to improve discriminative tasks, our proposed Denoising Model for Representation Learning of Person Re-Identification (DenoiseReID) is as below: Firstly, we utilize a well-trained backbone and keep it fixed throughout all subsequent procedures. This step is a free launch as we can easily use any publicly available backbone without requiring additional training time. This approach allows us to preserve the backbone’s inherent characteristics of semantic feature extraction. Given the backbone and an image, we can get a list of features. Next, we train denoising layers on those features. The weights of denoising layers are randomly initialized and their weights are not shared. The training process is the same as that in DDPM [20], where the only difference is that the denoising layer in DDPM takes a dynamic $t\in[1,T]$, and our denoising layers take fixed $n\in[1,N]$, where $n$ is the layer index, $T$ is denoise times and $N$ is backbone layer number as shown in Fig. 1(c). Finally, considering that the $N$ denoising layers consume additional execution latency, we propose a novel Feature Extraction and Feature Denoising Fusion Algorithm (FEFDFA). As shown in Fig. 1(d), FEFDFA merges parameters of extra denoising layers into weights of the existing embedding layers, thus enabling joint feature extraction and denoising without any extra computation cost. We also theoretically demonstrate the total equivalence before and after parameter fusion. Please see Section 3.3 and Eq (7) for more details.
@@
(1) We propose a novel Denoising Model for Representation Learning of Person Re-Identification (DenoiseReID), which innovatively integrates the denoising process, originating from generative tasks, into the discriminative tasks. It treats $N$ cascaded embedding layers of a backbone as $T$ times recursively proceeded denoising layers. This idea enables joint feature extraction and denoising is a backbone, thus making features more discriminative.
(2) We propose a novel Feature Extraction and Feature Denoising Fusion Algorithm (FEFDFA), which fuses the parameters of the denoising layers into the parameters of the corresponding embedding layers and theoretically demonstrates their equivalence. This contributes to a computation-efficient algorithm, which takes no extra latency.
(3) Extensive experiments on 4 ReID datasets verified that our proposed DenoiseReID can effectively improve feature performance in a label-free manner and performs better in the case of label-argumented supervised training or introduction of additional training data. We also extend our proposed method to large-scale (ImageNet) and fine-grained (e.g. CUB200, Oxford-Pet, etc.) image classifications, showing its effectiveness.
@@ 2 Related Work
Generative models learn the distribution of inputs before estimating class probabilities. A generative model is a model that learns the data generation process by learning the probability distribution of the input data and generating new data samples. The generative models first estimate the conditional density of categories $P(x|y=k)$ and prior category probabilities $P(y=k)$ from the training data. The $P(x)$ is obtained by the full probability formula. So as to model the probability distribution of each type of data. Generative models can generate new samples by modelling data distribution. For example, Generative Adversarial Networks (GANs) [17, 39, 21] and Variational Autoencoders (VAEs) [22, 48, 62, 58] are both classic generative models that generate real samples by learning potential representations of data distributions, demonstrating excellent performance in data distribution modeling. Recent research has focused on using diffusion models for generative tasks. The diffusion model was first proposed by the article [46] in 2015, with the aim of eliminating Gaussian noise from continuous application to training images. The DDPM [20] proposed in 2020 have made the use of diffusion models for image generation mainstream. In addition to its powerful generation ability, the diffusion model also has good denoising ability through noise sampling, which can denoise noisy data and restore its original data distribution.
@@
Discriminative models learn condition datatribution, e.g. $P(y|x)$, where $x$ is input data and $y$ is task-relative features, classiification task [1, 2, 12] map data into its tags, retrieval task [33, 57] map data into a feature space where similar data should be near otherwise farawary, detection task [44, 25] map data into is space position and size. Person Re-Identification (ReID) is a fine-grained retrieval task that solves the challenging task of identifying individuals in different camera views. Due to its fine-grained nature, it has a more discriminative feature, so we chose it as our benchmark task. Existing ReID methods can be grouped into hand-crafted descriptors [32, 38, 59] incorporated with metric learning [23, 31, 64] and deep learning algorithms [53, 52, 51, 40, 16, 9]. State-of-the-art ReID models often leverage convolutional neural networks (CNNs) [26] to capture intricate spatial relationships and hierarchical features within person images. Attention mechanisms [49, 14], spatial-temporal modeling [29, 28], and domain adaptation techniques [8] have further enhanced the adaptability of ReID models to diverse and challenging real-world scenarios.
Reparameterization of a model adopts an approximate equivalent new method to replace an existing modeling method. Reparameterization can reduce the number of parameters, reduce model computational complexity, and improve design flexibility. As RepVGG [13] proposed, a multi-branch structure is set up in the VGG network, which only has linear operations inside. During the inference stage, the parameters of multiple branches are merged into parameters of a single branch through reparameterization, without the need for retraining, thereby reducing the number of parameters and the time complexity of inference.
3 Denoising Representation for Person Re-Identification
3.1 Review Representation Learning of Person Re-Identification
The standard ReID pipeline consists of a backbone to learn representation followed by a classification loss and triplet loss for learning an embedding space where images of the same identity have high similarity. Common backbones include CNN series  [18] and Transformer series  [14].
In this work, we chose TransReID-SSL [37] as the baseline, which uses ViT as the backbone for feature extraction. As shown in the structural diagram on Fig. 2, the input image is first passed through PatchEmbed, which splits the image into chunks, and these chunks are fed to the transformer encoder. After that, it goes through $N$ Blocks for feature extraction and outputs a feature vector to represent this image.
3.2 Feature Extraction and Feature Denoising Unified Framework
In representing learning, feature representation is the most important factor that affects the person re-identification (ReID) accuracy, so we would like to learn features with better representation to improve the retrieval performance. We suppose that in the inference stage, the features obtained by backbone extraction are noisy, which may come from the input image or may be generated during the feature extraction process.
@@ 3.2 Feature Extraction and Feature Denoising Unified Framework
To solve the problem above, we propose a novel Feature Extraction and Feature Denoising Unified Framework (FEFDUF), which unifies feature extraction and feature denoising in a single backbone to obtain semantic and discriminative features. We refer to the diffusion modeling approach to denoise the noisy features through T-steps to obtain clean features. At the beginning, we use the features output from the backbone network as data samples for diffusion training, and get the noisy samples by continuously adding noise and learning through the network in order to simulate the data distribution of its features.
@@
where $X_{t}$ represents the feature vector output by the backbone in the inference stage, which we suppose contains some noise. $T$ is the denoising step size, representing the magnitude of the noise. We adjust $t$ appropriately based on different datasets and backbones to obtain the optimal denoising amplitude. In ViT-base and ViT-small backbone, $t$ is set to 10 and 15, respectively. According to $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$ denoise it step by step, and finally obtains $X_{0}$, which represents the clean feature after denoising.
Figure 2: Pipeline of our proposed method. ViT consists of $N$ cascaded transformer encoder layers. During the training phase (see the right side “Train Only” process), we freeze the backbone parameters and only train the extra denoising layers. In the inference stage (see the left side “Infer Only” process), we merge the parameters of denoising layers to corresponding encoder layers with the proposed Feature Extraction and Feature Denoising Fusion Algorithm (FEFDFA). So there is no extra inference latency cost. Please find definitions of $W$, $b$, $W_{D}$, $W^{\prime}$, $i$ and $b^{\prime}$ in Algorithm 2.
3.3 Feature Extraction and Feature Denoising Fusion Algorithm
As described in section 3.2, the proposed Feature Extraction and Feature Denoising Unified Framework (FEFDUF) could effectively polish noisy features but extra inference latency is introduced caused by recursive calling of the denoising layers.
To solve the problem above, we proposed a novel Feature Extraction and Feature Denoising Fusion Algorithm (FEFDFA) to fuse parameters of the extra denoising layers into parameters of existing backbone layers and get a computation-free method in the final. FEFDFA expands the linear layer in the MLP of each block into two branches, one for its original $FC$ module and the other for FEFDFA proposed in this paper. As shown on Fig. 2, during the training phase, we freeze the original parameters and only trained the FEFDFA. The training method is consistent with section 3.2, and the features are diffused and fed into the predictive noise network of FEFDFA for prediction. For details, please refer to Algorithm. 1. In the inference stage, we use the reparameterization method to combine the pre-trained $W$ parameters with the denoising module parameters, merging the two branches into a single branch without additional inference time. The derivation of parameter merging is as follows:
@@ 3.3 Feature Extraction and Feature Denoising Fusion Algorithm
Algorithm 1 Denoising Model Training
1:The number of feature layers in the backbone N, Features extracted from each layer $\{F_{i}\}_{i=1}^{N}$, The denoising module that needs to be trained $\{D_{i}(\cdot)\}_{i=1}^{N}$
@@
5:         $\epsilon\sim N(0,I)$: Randomly sample a Gaussian noise
6:         $X_{t}=\sqrt{\bar{a_{t}}}F_{i}+\sqrt{1-\bar{a_{t}}}\epsilon$: Forward diffusion process in Eq.(2)
@@
The FEFDFA we propose is based on feature level denoising and can be migrated to various downstream tasks. It denoises the features on each layer for better removal of noise at each stage, as the noise in the inference stage comes from multiple sources, which could be noise in the input image or noise generated while passing through the network. Denoising each layer avoids noise accumulation and gives better quality output. And according to the noise challenges brought by data in different scenarios, the denoising intensity can be adjusted by controlling $t$, $\beta_{t}$, and the number of denoising times, which has good generalization ability.
Algorithm 2 Denoising Model Inference
@@
6:     $F^{t-1}=W^{\prime}F^{t}+b^{\prime}$: Feature Extraction and Feature Denoising Fusion.
@@ 3.4 Unsupervised Learning Manner
Our proposed FEFDFA is suitable for label-free unsupervised training because its essence is a generative model that models data by learning its distribution. Thus the training Loss contains only the $Loss_{p}$ of prediction noise module:
$Loss_{p}=\left\|\epsilon-D_{\theta}(X_{t},t)\right\|$ (11)
where $\epsilon$ denotes the sampled noise, $X_{t}$ denotes the noise sample, $t$ denotes the diffusion step, and $D_{\theta}(X_{t},t)$ denotes the noise predicted by the model.
However, it is worth noting that our method is complementary to label if the label is available. $Loss_{ReID}$ is the original search loss for baseline, $\lambda$ is the trade-off parameter between two loss functions. The label-argumented learning is defined as:
$Loss=(1-\lambda)Loss_{ReID}+\lambda Loss_{p}$ (12)
Results in experiments section 4.1 shows the improved performance by introducing label information for supervised training.

### v2

# DenoiseRep: Denoising Model for Representation Learning

@@ Abstract
The denoising model has been proven a powerful generative model but has little exploration of discriminative tasks. Representation learning is important in discriminative tasks, which is defined as "learning representations (or features) of the data that make it easier to extract useful information when building classifiers or other predictors" [4]. In this paper, we propose a novel Denoising Model for Representation Learning (DenoiseRep) to improve feature discrimination with joint feature extraction and denoising. DenoiseRep views each embedding layer in a backbone as a denoising layer, processing the cascaded embedding layers as if we are recursively denoise features step-by-step. This unifies the frameworks of feature extraction and denoising, where the former progressively embeds features from low-level to high-level, and the latter recursively denoises features step-by-step. After that, DenoiseRep fus es the parameters of feature extraction and denoising layers, and theoretically demonstrates its equivalence before and after the fusion, thus making feature denoising computation-free. DenoiseRep is a label-free algorithm that incrementally improves features but also complementary to the label if available. Experimental results on various discriminative vision tasks, including re-identification (Market-1501, DukeMTMC-reID, MSMT17, CUHK-03, vehicleID), image classification (ImageNet, UB200, Oxford-Pet, Flowers), object detection (COCO), image segmentation (ADE20K) show stability and impressive improvements. We also validate its effectiveness on the CNN (ResNet) and Transformer (ViT, Swin, Vmamda) architectures. Code is available at https://github.com/wangguanan/DenoiseRep.
@@ 1 Introduction
Denoising Diffusion Probabilistic Models (DDPM)  [20] or Diffusion Model for short have been proven to be a powerful generative model [5]. Generative models can generate vivid samples (such as images, audio and video) by modeling the joint distribution of the data $P(X,Y)$, where $X$ is the sample and $Y$ is the condition. Diffusion models achieve this goal by adding Gaussian noise to the data and training a denoising model of inversion to predict the noise. Diffusion models can generate multi-formity and rich samples, such as Stable diffusion [45], DALL [43] series and Midjourney, these powerful image generation models, which are essentially diffusion models.
However, its application to discriminative models has not been extensively explored. Different from generative models, discriminative models predict data labels by modeling the marginal distribution of the data $P(Y|X)$. $Y$ can be various labels, such as image tags for classification, object boxes for detection, and pixel tags for segmentation. Currently, there are several methods based on diffusion models implemented in specific fields. For example, DiffusionDet [6] is a new object detection framework that models object detection as a denoising diffusion process from noise boxes to object boxes. It describes object detection as a generative denoising process and performs well compared to previous mature object detectors. DiffSeg [47] for image segmentation, which is a method of unsupervised zero-shot sample segmentation using pre-trained models (stable diffusion). It introduces a simple and effective iterative merging process to measure the attention maps between KL divergence and merge them into an effective segmentation mask. The proposed method does not require any training or language dependency to extract the quality segmentation of any image.
Figure 1: A brief description of our idea. (a) A typical denoising model for generative tasks recursively applies a denoising layer. (b) A naive idea that applies a denoising strategy to a discriminative model is applying a recursive denoise layer on the feature of a backbone and taking extra inference latency. (c,d) Our DenoiseRep first unifies the frameworks of feature extraction and denoising in a backbone pipeline, then merges parameters of the denoising layers into embedding layers, making the feature more discriminative without extra latency cost.
@@
To resolve all the difficulties above and efficiently apply the denoising process to improve discriminative tasks, our proposed Denoising Model for Representation Learning (DenoiseRep) is as below: Firstly, we utilize a well-trained backbone and keep it fixed throughout all subsequent procedures. This step is a free launch as we can easily use any publicly available backbone without requiring additional training time. This approach allows us to preserve the backbone’s inherent characteristics of semantic feature extraction. Given the backbone and an image, we can get a list of features. Next, we train denoising layers on those features. The weights of denoising layers are randomly initialized and their weights are not shared. The training process is the same as that in DDPM [20], where the only difference is that the denoising layer in DDPM takes a dynamic $t\in[1,T]$, and our denoising layers take fixed $n\in[1,N]$, where $n$ is the layer index, $T$ is denoise times and $N$ is backbone layer number as shown in Fig. 1(c). Finally, considering that the $N$ denoising layers consume additional execution latency, we propose a novel feature extraction and feature denoising fusion algorithm. As shown in Fig. 1(d), the algorithm merges parameters of extra denoising layers into weights of the existing embedding layers, thus enabling joint feature extraction and denoising without any extra computation cost. We also theoretically demonstrate the total equivalence before and after parameter fusion. Please see Section 3.3 and Eq (7) for more details.
@@
(1) We propose a novel Denoising Model for Representation Learning (DenoiseRep), which innovatively integrates the denoising process, originating from generative tasks, into the discriminative tasks. It treats $N$ cascaded embedding layers of a backbone as $T$ times recursively proceeded denoising layers. This idea enables joint feature extraction and denoising is a backbone, thus making features more discriminative.
(2) The proposed DenoiseRep fuses the parameters of the denoising layers into the parameters of the corresponding embedding layers and theoretically demonstrates their equivalence. This contributes to a computation-efficient algorithm, which takes no extra latency.
(3) Extensive experiments on 4 ReID datasets verified that our proposed DenoiseRep can effectively improve feature performance in a label-free manner and performs better in the case of label-argumented supervised training or introduction of additional training data. We also extend DenoiseRep to large-scale (ImageNet), fine-grained (CUB200, Oxford-Pet, Flowers) image classifications, object detection (COCO) and image segmentation (ADE20K), showing its scalability.
@@ 2 Related Work
Generative models learn the distribution of inputs before estimating class probabilities. A generative model learns the data generation process by learning the probability distribution of the input data and generating new data samples. The generative models first estimate the conditional density of categories $P(x|y=k)$ and prior category probabilities $P(y=k)$ from the training data. The $P(x)$ is obtained by the full probability formula. So as to model the probability distribution of each type of data. Generative models can generate new samples by modelling data distribution. For example, Generative Adversarial Networks (GANs) [17, 39, 21] and Variational Autoencoders (VAEs) [22, 48, 62, 58] are both classic generative models that generate real samples by learning potential representations of data distributions, demonstrating excellent performance in data distribution modeling. Recent research has focused on using diffusion models for generative tasks. The diffusion model was first proposed by the article [46] in 2015, with the aim of eliminating Gaussian noise from continuous application to training images. The DDPM [20] proposed in 2020 have made the use of diffusion models for image generation mainstream. In addition to its powerful generation ability, the diffusion model also has good denoising ability through noise sampling, which can denoise noisy data and restore its original data distribution.
@@
Discriminative models learn condition distribution, i.e. $P(y|x)$, where $x$ is data and $y$ is task-specific features. For example, classification tasks [1, 2, 12] map data to tags, retrieval tasks [33, 57] map data to a feature space where similar data should be near otherwise faraway, detection task [44, 25] map data to space position and size. Person Re-Identification (ReID) is a fine-grained retrieval task which identifies individuals among disjoint camera views. Considering its challenge to feature discrimination, we take ReID as the major benchmark task and the others as auxiliary benchmarks. Existing ReID methods can be grouped into hand-crafted descriptors [32, 38, 59] incorporated with metric learning [23, 31, 64] and deep learning algorithms [53, 52, 51, 40, 16, 9]. State-of-the-art ReID models often leverage convolutional neural networks (CNNs) [26] to capture intricate spatial relationships and hierarchical features within person images. Attention mechanisms [49, 14], spatial-temporal modeling [29, 28], and domain adaptation techniques [8] have further enhanced the adaptability of ReID models to diverse and challenging real-world scenarios.
3 DenoiseRep: Denoising Model for Representation Learning
3.1 Review Representation Learning
Representation learning plays a pivotal role in discriminative tasks, which is defined as "learning representations (or features) of the data that make it easier to extract useful information when building classifiers or other predictors" [4]. A common architecture of discriminative tasks consists of a vision backbone to extract discriminative features (e.g., ResNet [18], ViT [14]) and a task-specific head that operates on these features (e.g. MLP [24] for classification, RCNN [44] for object detection, FCN [long2015fully] for segmentation). It is evident that the vision backbone is central to representation learning. In this paper, we introduce a novel Denoising Model for Representation Learning (DenoiseRep), which integrates feature extraction and feature denoising within a single vision backbone. This approach aims to enhance the discriminative power of the features extracted.
3.2 Joint Feature Extraction and Feature Denoising
@@
We refer to the diffusion modeling approach to denoise the noisy features through T-steps to obtain clean features. At the beginning, we use the features output from the backbone network as data samples for diffusion training, and get the noisy samples by continuously adding noise and learning through the network in order to simulate the data distribution of its features.
@@
where $X_{t}$ represents the feature vector output by the backbone in the inference stage. $T$ is the denoising step size, representing the magnitude of the noise. We adjust $t$ appropriately based on different datasets and backbones to obtain the optimal denoising amplitude. According to $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$ denoise it step by step, and finally obtains $X_{0}$, which represents the clean feature after denoising.
Figure 2: Pipeline of our proposed DenoiseRep. ViT consists of $N$ cascaded transformer encoder layers. During the training phase (see the right side “Train Only” process), we freeze the backbone parameters and only train the extra denoising layers. In the inference stage (see the left side “Infer Only” process), we merge the parameters of denoising layers to corresponding encoder layers. So there is no extra inference latency cost. Please find definitions of $W$, $b$, $W_{D}$, $W^{\prime}$, $i$ and $b^{\prime}$ in Algorithm 2.
3.3 Fuse Feature Extraction and Feature Denoising
As described in Section 3.2, the proposed method above could effectively improve the discriminability of features. Still, extra inference latency is introduced caused by recursive calling of the denoising layers. To solve the problem, we propose to fuse parameters of feature denoising layers into parameters of existing embedding layers of the backbone. The core idea is to expand the linear layer each transformer encoder block into two branches, one for its original embedding layer and the other for extra denoising layer. As shown in Fig. 2, during the training phase, we freeze the original embedding layers and only train the denoising layers. The training method is consistent with section 3.2, and the features are diffused and fed into the denoising layers. Please refer to Algorithm 1 for more details. In the inference stage, we fuse the pre-trained parameters of embedding and denoising layers, merging the two branches into a single branch without additional inference time. Please note that, here we take the transformer architecture as an example, but DenoiseRep is suitable for CNN architecture. We demonstrate its scalibity on CNN in experiments. The derivation of parameter merging is as follows:
@@ 3.3 Fuse Feature Extraction and Feature Denoising
Algorithm 1 Training
1:The number of feature layers in the backbone N, features extracted from each layer $\{F_{i}\}_{i=1}^{N}$, the denoising module that needs to be trained $\{D_{i}(\cdot)\}_{i=1}^{N}$.
@@
5:         $\epsilon\sim N(0,I)$: Randomly sample a Gaussian noise.
6:         $X_{t}=\sqrt{\bar{a_{t}}}F_{i}+\sqrt{1-\bar{a_{t}}}\epsilon$: Forward diffusion process in Eq.(2).
@@
Our proposed DenoiseRep is based on feature-level denoising and can be migrated to various downstream tasks. It denoises the features on each layer for better removal of noise at each stage, as the noise in the inference stage comes from multiple sources, which could be noise in the input image or noise generated while passing through the network. Denoising each layer avoids noise accumulation and gives better quality output. And according to the noise challenges brought by data in different scenarios, the denoising intensity can be adjusted by controlling $t$, $\beta_{t}$, and the number of denoising times, which has good generalization ability.
Algorithm 2 Inference
@@
6:     $F^{t-1}=W^{\prime}F^{t}+b^{\prime}$: Fuse feature extraction and feature denoising.
@@ 3.4 Unsupervised Learning Manner
Our proposed DenoiseRep is label-free because its essence is a generative model that models data by learning its distribution. Thus the training loss contains only the $Loss_{p}$ of denoising layers:
$Loss_{p}=\sum_{i=1}^{N}\left|\epsilon_{i}-D_{\theta_{i}}(X_{t_{i}},t_{i})\right|$ (11)
where $\epsilon$ denotes the sampled noise, $N$ denotes the number of denoising layers, $X_{t}$ denotes the noise sample, $t$ denotes the diffusion step, and $D_{\theta}(X_{t},t)$ denotes the noise predicted by the denoising layer.
However, it is worth noting that our method is complementary to label if the label is available. $Loss_{l}$ is the task-specific supervised loss with label, $\lambda$ is the trade-off parameter between two losses. The label-argumented learning is defined as:
$Loss=(1-\lambda)Loss_{l}+\lambda Loss_{p}$ (12)
Results in experiments Section 4.1 shows the improvement from label.